OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (5, 7) and (8, 22)? y = −5x + 18 y = 5x − 18 y = −5x − 18 y = 5x + 18

OpenStudy (jack1):

y = mx + c is slpoe intercept form... yeah?

OpenStudy (anonymous):

how would i go about finding that @Jack1

OpenStudy (jack1):

(5, 7) and (8, 22) are your points on the line, so lets call them (x1, y1) and (x2,y2)

OpenStudy (anonymous):

ok

OpenStudy (jack1):

gradient (m) = rise / run = (y2 - y1)/(x2-x1) so sub in those values and tell me what you get

OpenStudy (jack1):

@aleisha96

OpenStudy (anonymous):

is it 22-7/8-5?

OpenStudy (jack1):

perfect, so that = 15/3 = 5/1 = 5 so m = 5

OpenStudy (jack1):

so now u know the answer cant be a or c, as their gradient is -5, and ours is 5 so must be b or d

OpenStudy (anonymous):

so how would we go about finding x and c?

OpenStudy (jack1):

x is a variable, so is y but c and m are fixed y = mx + c y = 5x + c now take one of the points, lets use (5, 7) (x, y) so sub those values for x and y into the below equation, and solve for c y = 5x + c

OpenStudy (anonymous):

isnt slope intercept formula y=mx+b

OpenStudy (jack1):

ok, same thing though y = 5x + b sub in and solve for b

OpenStudy (anonymous):

so its 5x+ ????

OpenStudy (jack1):

yeah... its y = 5x + b and a point on that line is (5,7) this point is an x,y coordinate so sub the x value of 5 into the above exuationm and sub the y value of 7 into the smae equation, and what are you left with>

OpenStudy (anonymous):

5x+7?

OpenStudy (jack1):

no y = 5x + b sub in the y value 7 = 5x + b now sub in the x value 7 = 5(5) + b

OpenStudy (jack1):

so expand it out and leave the b on one side of the equals sign alone 7 = 5(5) + b 7 = 25 + b therefore b = ...?

OpenStudy (anonymous):

-18

OpenStudy (jack1):

perfect,

OpenStudy (anonymous):

so its b!

OpenStudy (jack1):

yep! ;D

OpenStudy (anonymous):

can you help me with another question

OpenStudy (jack1):

of course, shoot

OpenStudy (anonymous):

Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3. −15 −8 10 25

OpenStudy (jack1):

f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3. a simple way to look at this is g(x) = x − 7... lets call g(x) "p" instead p = x − 7 now f(x) = 2x + 5, and when f(p)

OpenStudy (anonymous):

ok

OpenStudy (jack1):

f(x) = 2x + 5 f(p) = 2p + 5 ... and remember p = x − 7, so f(p) = 2(x − 7) + 5 f(p) = 2x − 14 + 5 f(p) = 2x − 9 now the final part is f(g(x)) when x = −3 f(p) = 2x − 9

OpenStudy (jack1):

now the final part is f(g(x)) when x = −3 f(p) = 2(-3) − 9 f(p) = -6 − 9 f(p) = -15 so f(g(x)) at f(g(-3)) is -15

OpenStudy (jack1):

does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

sweet. nite

OpenStudy (anonymous):

@jack 1 thanks